# Evaluating the Defense Potential of Machine Unlearning against Membership Inference Attacks

### v3

@@ 6.5 Sensitivity Analysis
The sensitivity analysis of the model’s performance across various learning rates (see Fig 27) reveals a key trade-off between model accuracy stability and resilience to Membership Inference Attacks (MIA). The first chart demonstrates that, across all learning rates, the model achieves high accuracy for both forgotten and retained data, with accuracies consistently ranging from 0.97 to 1.0 for forgotten data and slightly lower but still stable accuracy for retained data. Lower learning rates, such as (0.0005), show minimal fluctuations in performance, indicating greater stability over time. Higher learning rates, while still producing high accuracy, introduce minor variability, suggesting that these may lead to faster learning at the cost of small performance fluctuations.
In contrast, the second chart (see Fig 28), focusing on MIA accuracy, highlights that higher learning rates significantly improve the model’s resistance to MIA attacks. The MIA accuracy for forgotten data declines more rapidly with higher learning rates, suggesting that these configurations reduce the model’s vulnerability to attacks more effectively. Conversely, lower learning rates, although stable in terms of accuracy, leave the model more susceptible to privacy risks, particularly for retained data where MIA accuracy remains higher. This analysis suggests that while lower learning rates offer more consistent accuracy, higher learning rates provide better protection against MIA attacks, especially for forgotten data, striking a crucial balance between stability and privacy.
Figure 27: sensitivity mia accuracy
Figure 28: sensitivity mia accuracy
@@ 7 Conclusion
This paper investigates the effectiveness of Machine Unlearning as a defense mechanism against MIAs across a range of datasets and model architectures. Our experimental results demonstrated that Machine Unlearning can indeed play a significant role in reducing a model’s vulnerability to MIAs, though the effectiveness of this defense is highly dependent on the choice of unlearning algorithm.
Among the algorithms evaluated, i.e., NegGrad, SCRUB and SFTC, the latter consistently outperformed the others in terms of both unlearning effectiveness and preservation of model utility. It achieved a successful balance by thoroughly erasing the influence of forget data while maintaining high accuracy on retained data. This aspect is particular important in privacy-sensitive applications, where compliance with regulations like the GDPR requires robust data removal, but without compromising the model’s predictive power.
NegGrad, although effective in unlearning the forget set, exhibited significant overfitting, leading to lower test accuracy and greater degradation in model performance. SCRUB, on the other hand, offered moderate success, preserving model utility but being less effective in fully forgetting the targeted data, as evidenced by relatively stable forget accuracy in several datasets. Both methods showed weaknesses when compared to SFTC, particularly in terms of maintaining generalization and reducing MIA risks.
The sensitivity analysis revealed that higher learning rates generally improved the effectiveness of unlearning, with SFTC proving less sensitive to learning rate changes than NegGrad or SCRUB. This robustness makes SFTC more adaptable and practical for real-world ML applications, where hyperparameter tuning may be limited.
Despite the promise of Machine Unlearning, it is clear that this technique is not a universal safeguard against MIAs. While SFTC shows considerable potential in mitigating the risks, the choice of unlearning algorithm and its interaction with specific datasets and architectures are critical factors in determining the overall level of protection. Moreover, residual vulnerabilities may still exist in models post-unlearning, as attackers could exploit subtle traces of the forget data.

### v4

@@ 7.4 Sensitivity Analysis
(a) Neg Grad at cifar with lr 0.00001.
(b) Neg Grad at cifar with lr 0.0001.
(c) Neg Grad at cifar with lr 0.00001.
Figure 12: The unlearning curve through MIA for three different learning rates of Neg Grad for the cifar dataset and the architecture 1.
Throughout the implementation of our experiments, we observed a profound sensitivity regarding the learning rate of the unlearning algorithms. Minor adjustments to this hyperparameter resulted in substantial shifts in both the effectiveness of the MIA and the overall success of the unlearning process (see Figure 12). For instance, with a learning rate of 0.00001, the MIA Forget Accuracy reached 0.830, surpassing the Retain accuracy, which indicates a highly effective attack on the forget set. Consequently, we conducted extensive grid searches across various learning rates, ultimately selecting the values that provided the most robust defense against MIAs. This high sensitivity underscores a critical challenge in machine unlearning that warrants rigorous exploration in future research.
@@ 8 Conclusion
This paper investigates the efficacy of Machine Unlearning as a defense mechanism against Membership Inference Attacks (MIAs) across a diverse range of datasets and model architectures. Our experimental results demonstrate that Machine Unlearning can effectively mitigate a model’s vulnerability to MIAs for the forget set, but this process requires a very careful consideration of the parameters. However, the success of this defense is critically contingent upon several factors: the selection of the unlearning algorithm, hyperparameter tuning (specifically the learning rate), the initial effectiveness of the MIA, the degree of model overfitting, and the number of unlearning epochs utilized.
In real-world scenarios, the application of 30 unlearning epochs typically results in the predictive collapse of the model. In our experiments, we deliberately extended the unlearning process to this duration to examine the unlearning dynamics through the perspective of MIA susceptibility.
The experimental results across the Negative Gradient, Scrub, and SFTC algorithms reveal that the effectiveness of machine unlearning as a defense against Membership Inference Attacks (MIA) is deeply tied to both model complexity and the specific unlearning mechanism utilized. While the Negative Gradient method provides a baseline level of protection, it often results in generalized degradation where the membership signals of both forget and retain sets decline simultaneously, particularly in datasets like Texas-100 and Purchase-100. The Scrub algorithm offers a more aggressive approach but is prone to significant instability in simple architectures and can paradoxically increase model transparency in certain scenarios, creating a “bell-curve” risk profile where data becomes more identifiable before being masked.
The SFTC algorithm is characterized by a distinct “divergence” effect, where the MIA Forget Accuracy decreases while the MIA Retain Accuracy is consolidated or even strengthened. This phenomenon is particularly evident in high-capacity models. The internal mechanics of SFTC involve intensive fine-tuning on the retain data alongside the selective “confusion” of the forget data. Because the student model is trained rigorously to mirror the teacher’s high confidence on retain samples, it inadvertently reinforces their membership signals, making them more identifiable to an adversary. And that is obvious in our experiments where the MIA Retain accuracy is increasing.
In contrast, the SCRUB algorithm also employs a dual-phase approach: a maximize phase to induce forgetting and a minimize phase for fine-tuning on the retain data. However, our experiments show that the resulting increase in MIA Retain Accuracy is generally negligible compared to SFTC. Instead, SCRUB focuses on maintaining a significant drop in MIA Forget Accuracy. While SCRUB uses fine-tuning to recover model utility, it typically aims for stabilization rather than the aggressive reinforcement of the retain set’s signal seen in SFTC.
Despite the promise of Machine Unlearning, it is clear that this technique is not a universal safeguard against MIAs. Residual vulnerabilities may still exist in models post-unlearning, as attackers could exploit subtle traces of the forget data.
@@
Acknowledgements.
We thank the anonymous reviewers for their insightful and constructive comments, which have significantly improved the quality of this work.
Declarations
The authors declare that no funds, grants, or other support were received during the preparation of this manuscript. The authors have no competing interests to declare.